Question 1 (20 marks)

Brilliant Company uses the periodic inventory system. The company’s beginning inventory and later purchases of inventory during June 2020 were as below:

Jun 01: Beginning inventory, 30 units @ $22 per unit.

Jun 17: Inventory purchased, 95 units @ $25 per unit.

Jun 25: Inventory purchased, 50 units @ $28 per unit.

Brilliant company sold 110 units during the month of June 2020.

Required:

  1. Calculate the cost of goods sold and ending inventory for June 2020 using both the Weighted Average cost and First in, first out (FIFO) method. (14 marks)

  1. Also record the journal entries for the June 17 purchase of inventory and when 80 units from this lot were sold for $40 each. Assume both purchase of inventory and its sale are on account. (6 marks)
Answer #1

If you like my explanation, please give a "thumbs up"

A B с D E F G H K 1 price Cost w N A 22 660 units 1-Jun 17-Jun 25-Jun 30 95 50 175 25 2375 28 1400 4 5 75 4435 6 7 110 8 Sale

For calculation ref:

A B с D E F G H 1 J units 1 2 43983 30 3 95 43999 44007 price Cost 22 =D2 C2 25 =D3*C3 28 =D4*C4 =SUM(D2:04) =SUM(E2:E4) 4 50